The molarity of strandium hydroxide is 8.834*10^-2
is then used to titrate 26.81 mL of a solution of bromoacetic acie (BrH,CCo0H), An endpoint is reached when 20.59 mL of strontum h hydroxide has been added 2. The Determine the molarity (in mol/L) of the bromoacetic acid solution, Report your answer to four significant figures Submit Answer Tries 0/3
Answer #1

Sr(OH)2 (aq.) + 2 BrCH2COOH (aq.) ------------> Sr(BrCH2COO)2 (aq.) + 2 H2O (l)

Useful formula,

M1 * V1 / n1 = M2 * V2 / n2

8.834 * 10-2 * 20.59 / 1 = M2 * 26.81 / 2

M2 = 0.136 M

Molarity of bromoacetic acid = 0.136 mol / L